MATH 136 – Advanced Placement Calculus Problem Set Guidelines Fall 2009

Weekly homework assignments will be posted each Friday on the course homepage and collected the following Friday. These assignments will usually consist of 20-30 problems (depending on level of difficulty, how many parts there are to each question, and so forth). Because of the size of the class, it will usually be the case that only a selection of the problems will be graded – this is to ensure that you get timely feedback. Here are some guidelines to follow while doing your assignments, for our mutual benefit:

• Please be neat! If your first try on a problem needs to be erased, scratched out, heavily edited, etc. then recopy the problem before turning it in. I reserve the right to refuse to correct papers that are too messy to deal with in a reasonable way. If that happens, I will give you the option of recopying and handing in the assignment at the next class.

• Write up the problems in order. Use only one side of the page (don’t worry about saving paper), and staple your assignment together, with the pages in the proper order.

• Please do not put several solutions side by side on the page – that makes it more difficult and time-consuming to find all the problems and give you the credit you are due.

• Leave space for comments.

• Show enough detail so that a confused student in the class could follow your solution. Don’t expect that the reader already knows how to solve the problem. Just writing an answer is never sufficient.

• Start the assignment early (i.e. the day it is posted). This will allow you ample time to consult with me, the Calculus Workshop, or your classmates if you get stuck on some problems. If you start the assignment on the day before it is due, I can promise you that you will be very unhappy!

• If you work with other students to solve a problem, you must write up your final solutions independently. Plan on taking scrap notes when you work out the problem first, and then write up your solution in a clean form in your own words, by yourself afterwards. Add a note to your solution listing the other people you consulted.